Grade 12Mechanics

A binary system consists of two stars of equal mass which revolve in circular orbit around their center of mass.If the orbital speed of each star is 220Km/s and theorbital period of each is 14.4 days, the mass of each star is ___________

Akshay Krishan.S.R
8 Years agoGrade 12

1 Answer

Khimraj
8 Years ago
For binary star system two force act on a star and in equilibrium condition
mv2/r = Gmm/(2r)2
So m = 4v2r/G.
and T = 2\Pir/v.
So r = Tv/2\Pi
So m = 2Tv3/\PiG.
T = 14.4 days = 14.4*24*60*60 secons
v = 220km/s = 220000m/s.
Now put these value in formula of mass. Then you will get your answer.